[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Jul 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Jul 25 19:09:43 UTC 2017


HSU Shift Log for Tue Jul 25 2017 15:09:37 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: JTWC called and said they were having trouble getting the  
forecasts for the EPac storms into their ATCF. Monica was notified. Problem  
could have resolved itself.

N-AWIPS issues: GOES16 imagery stopped coming in after 1441Z until 1505Z--  
then it all came in at once. Talked to Mello and he coordinated with Monica  
to fix problem. Then it stopped again at 1519Z. Seems to be coming in late  
in big chunks. Disk filling up?

AWIPS-2/NCP issues:

Other Techincal issues: FSSE guidance arrived very late today, after 1415Z  
for Hilary. It isn't very usable if we get it that late.

Miscellaneous issues: G16 imagery looks amazing in AWIP2 - it is a big  
upgrade over NAWIPS especially for IR.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170725/eef662f9/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list